OpenStudy (anonymous):

Write the point-slope equation of the line through (-2, 1) with a slope of 3

OpenStudy (anonymous):

the equation for point slope is y-y1=m(x-x1) you can plug in the point you have and the slope by going... y-(1)=3(x-(-2)) and simplify, y-1 = 3(x+2)